write a quadratic equation in the variable x havin the given numbers as solutions type the equation in standrd form, ax^2+bx+c=0
solution 2, only solution

Answer by stanbon(75887)

write a quadratic equation in the variable x havin the given numbers as solutions type the equation in standrd form, ax^2+bx+c=0
solution 2, only solution
------------
f(x) = (x-2)^2
f(x) = x^2-4x+4
=======================
Cheers,
Stan H.